OpenStudy (anonymous):

Can someone explain to me, STEP BY STEP how to do this problem please, its Algebra 2-Complex numbers. -6(x+5)^2=120

hartnn (hartnn):

first divide both sides by -6 what u get ?

hartnn (hartnn):

and u should know that \(\huge \sqrt{-1}=i\)

OpenStudy (anonymous):

6(x+5)^2+120 = (x+5)^2 - 20i^2 =0 ( and then factorise ) as i^2 =-1

OpenStudy (anonymous):

hartn good explanation and asking for participation . well done !!

OpenStudy (anonymous):